# Multi-Mode Array Filtering of Resonance Fluorescence

### v2

@@ I Introduction
The majority of studies on atomic resonance fluorescence at this time were largely focused on the entire fluorescence spectrum of the atom. It was at the end of the 1970s when some of the first investigations were conducted on correlations between photons from different components of the Mollow triplet, as opposed to the whole spectrum [9, 10, 11, 12, 13], with the first experimental results reported by Aspect, Roger, Dalibard and Cohen-Tannoudji [14]. To achieve the separation of the two side-peaks, the atomic fluorescence was split into two channels, with each channel passing through an interference filter tuned to the respective target frequency, and modeled as a single-mode optical cavity [15, 16, 17].
With this increased interest in measuring photon correlations of frequency components came interest in the effect that spectral filtering had on measurements [18, 19, 20, 21], as well as interest in other methods of calculating and measuring these correlations. In 2012, del Valle et al. proposed a novel method for measuring frequency-resolved photon correlations [22], consisting of a system of interest weakly coupled to $N$ two-level “detector” atoms, where $N$ is the order of the correlation function of interest. This method has been used extensively in theoretical studies due to its relatively simple implementation [23, 24, 25, 26, 27, 26, 28, 29]. There have been other methods discussed over the years, some with more theoretical leaning, such as: signal processing methods [30]; perturbation approaches [31]; frequency resolved Monte-Carlo, or quantum trajectory methods [32, 33, 34]. Some other methods, such as the eigenvalue decomposition of Liouvillian superoperators developed by Kamide et al [35, 36], allow for the modeling of not only Lorentzian type filters, but for filters with any frequency response.
@@
We also desire that the filter has no effect on the evolution of the source system. This is usually achieved by assuming a vanishingly small coupling of the source system to the filter, such that any back-action can be neglected. Holdaway et al. [31] take this approximation a step further with an algebraic expansion of the source-filter system with respect to the coupling parameter. We can, however, ensure there is no back-action at all by cascading the output of the source system into the filter, using cascaded open quantum systems theory [37, 38].
@@
We begin in Section II by introducing the foundations of our filtering model, including a brief discussion of single-mode cavities and rectangular filters. We then present analytic and numerical results for the temporal and frequency response of the multi-mode array filter for different parameters. In Section III we review the dynamics and key phenomena of a resonantly driven two-level atom, which will serve as the source system with which to test the effectiveness of the filter. We then introduce the full atom-filter coupled system in Section IV and the moment equations relevant to solving for the two-time correlation functions. In Section V we present analytic results for frequency-filtered auto- and cross-correlations, comparing the multi-mode array-filtered results with those of a standard Lorentzian filter and approximate ideal solutions derived in the dressed-state basis. Finally, we conclude with a summary of key results and a brief discussion on possible experimental implementations.
@@ II.1 Single-mode Lorentzian filter
We begin by considering a standard frequency filtering device, modeled as a single-mode cavity. Applying a coherent driving, with driving strength $\mathcal{E}_{d}$, the Hamiltonian modeling this system is
@@ II.2 The rectangular filter
as shown in Figs. 1(a) and (b). The infinitely sharp frequency response of the sinc, or rectangular, filter, would allow us to maximally increase the bandwidth, and thus the temporal response, of the filter response, much more than a conventional Lorentzian filter, while still maintaining good frequency isolation. Unfortunately, perfect sinc filters are physically impossible to realize due to their non-causal nature, that is, the temporal response depends on future inputs. By neglecting the negative temporal response of the sinc function with $\theta(t)\mathrm{sinc}(at)$, where
@@
is the Heaviside step function, we can realize a causal sinc filter. We see in Fig. 1(d), however, that the frequency response, as calculated from the Fourier transform, changes drastically. The infinitely sharp sides of the rectangular response have been replaced with large slow-decaying wings, together with a large dip in the center of the response.
Fortunately we are able to recover some of the rectangular nature by introducing a slight phase modulation, or delay, to the temporal response, $\theta(t)\mathrm{sinc}(at-\pi)$ (Fig. 1(e)). With this slight delay, the Fourier transform, Fig. 1(f), shows a much sharper cut-off than the non-delayed response in Fig. 1(d). By increasing the delay even further, we can “recover” much more of the sinc function, resulting in an even more rectangular frequency response. Unfortunately, a greater modulation also results in a greater temporal delay, which is undesirable for measuring accurate photon correlations. We will see in the following section that we are able to recreate this behavior with a cavity based model.
@@ II.3 The multi-mode array filter
Figure 2: Schematic of a coherently driven multi-mode array filter mode. The input field is evenly split into each individual two-sided cavity mode, where a mode-dependent phase modulation is applied. To achieve this equal coupling of the input light, the input field is passed through an array of beam splitters with increasing intensity reflectivity coefficient.
@@
We then assume that the fields transmitted through the out-coupling mirrors of the cavities are incident on a single photon detector, such that the total electric field at the detector is given by
@@
where $\omega_{j}$ is the frequency of the $j^{\mathrm{th}}$ cavity mode, $\epsilon_{0}$ is the electric permittivity of free space, and $A$ is the quantization area of the field. With frequencies in the optical range ($\sim 10^{15}$ Hz), we assume that the effective halfwidth of the multi-mode array filter, $N\delta\omega$, is much smaller than the central mode frequency $\omega_{0}$. The total field may then be approximated as a sum over the annihilation operators:
@@
is the normalized collective mode annihilation operator. For the single-mode case, $N=0$, $\kappa$ is also the halfwidth of the filter response. We therefore define the “effective halfwidth” of the multi-mode array filter as
@@ II.3.1 Temporal response
Figure 3: Normalized field amplitude response to an impulse driving, Eq. (II.3.1), with $m=0$ (blue, solid), $m=1$ (orange, dashed), and $m=2$ (green, dash-dot). Also shown is the sinc term of Eq. (II.3.1) (grey, dotted), showing the effect that the cavity decay, $\kappa$, has on the temporal response. Other parameters are $N=20$ and $\left(\delta\omega,\kappa,\omega_{c}\right)=\left(0.1,0.2,0.0\right)|\mathcal{E}_{d}|.$
@@
We see, then, that this proposed model behaves as we expect from a sinc-response filter, i.e., the temporal response is a positive-sided sinc function, with halfwidth $N\delta\omega$ and a temporal delay of $m\pi$. There is, however, an extra term that arises due to individual cavity modes which make up the multi-mode array filter: the exponential decay and oscillatory term $e^{-\left(\kappa+i\omega_{c}\right)t}$. The oscillatory term, $e^{-i\omega_{c}t}$, simply centers the response in frequency space at $\omega_{c}$. The cavity decay term, however, can have a much more significant effect on the filter’s response. We demonstrate this in Fig. 3, where the temporal response for the phase-shifted multi-mode array (orange, dashed), with $m=0,1$, or 2, is plotted against the sinc term in Eq. (II.3.1). The cavity decay causes the collective temporal response to deviate from the pure sinc term; an effect which is much more noticeable for larger $\kappa$. We can mitigate this effect, and thus achieve a more ideal sinc response, by ensuring the individual cavity decay rate is much smaller than the effective halfwidth of the array filter, i.e., $\kappa\ll N\delta\omega$.
@@ II.3.2 Frequency response
Figure 4: Normalized frequency response for a single-mode cavity with $\kappa=|\mathcal{E}_{d}|$ (grey, dotted) and a multi-mode array filter with (a) no phase modulation $(m=0)$, from Eq. (24), (b) $\pi$-phase modulation $(m=1)$, and (c) $2\pi$-phase modulation $(m=2)$, with $K=|\mathcal{E}_{d}|$ (blue, solid), $K=2|\mathcal{E}_{d}|$ (orange, dashed), and $K=4|\mathcal{E}_{d}|$ (green, dot-dash). Other parameters are $N=80,\omega_{c}=0$, and, for the multi-mode array filter, $\kappa=0.07|\mathcal{E}_{d}|$.
@@
$|\bar{A}(\omega)|^{2}$ $=\frac{|\mathcal{E}_{d}|^{2}}{\delta\omega^{2}\left(2N+1\right)^{2}}$ (24)
@@
We plot the collective frequency response in Fig. 4 for three different sizes of phase modulation – (a) $m=0$, calculated from Eq. (24), (b) $m=1$ , and (c) $m=2$ – for three different effective halfwidths: $N\delta\omega/|\mathcal{E}_{d}=1$ (blue, solid), 2 (orange, dashed), and 4 (green, dash-dot). By setting $m=0$, the collective response of the multi-mode array filter produces the “bunny-ear” like response, characteristic of the Fourier transform of the positive-sided sinc function in Fig. 1(d). Applying a slight modulation, with $m=1$, we see that the collective response closely resembles that of Fig. 1(f), with much sharper frequency cut-offs than a standard Lorentzian (grey, dotted). We cannot, however, push the phase modulation too far. As previously mentioned, an increase in the size of the modulation also increases the temporal response of the filter. We also see a noticeable dip in the center of the frequency response when $m=2$, in Fig. 4(c).
@@ III.2 The Mollow triplet
is the normalized steady-state first-order correlation function for the driven two-level atom. In the steady state there exist quantum fluctuations, describable with operators $\Delta\sigma_{\pm}=\sigma_{\pm}-\langle\sigma_{\pm}\rangle_{ss}$, and so we decompose the fluorescence spectrum into coherent and incoherent components:
@@
In the weak driving regime, $\Omega\ll\gamma$, the incoherent power spectrum of the atom is Lorentzian, centered at the atomic frequency $\omega_{A}$, and with halfwidth $\gamma/2$. As the Rabi frequency increases, and thus incoherent scattering dominates, side-peaks emerge at $\omega_{A}\pm\Omega$, giving shape to the Mollow triplet [2, 2], as shown in Fig. 5.
@@ III.3 Atomic dressed states
Each of the Mollow triplet peaks can be attributed to transitions amongst the atomic dressed states [42, 43] – eigenstates of the combined atomic and quantized field Hamiltonians. By diagonalizing the Hamiltonian, Eq. (26), we can express the eigenstates in terms of the atomic bare states as:
@@
These dressed states form an energy level ladder, where each pair are separated by an energy of $\hbar\omega_{A}$ in the quantized radiation field. In the strong driving regime, where the two dressed states are well separated, three sets of dressed state transitions occur, as depicted in Fig. 6: |$\ket{d}\rightarrow\ket{u}$ with frequency $\omega_{A}-\Omega$, |$\ket{u}\rightarrow\ket{u}$ and |$\ket{d}\rightarrow\ket{d}$ with frequency $\omega_{A}$, and |$\ket{u}\rightarrow\ket{d}$ with frequency $\omega_{A}+\Omega$. We can express the atomic operators in terms of these dressed states, allowing us to rewrite the atomic master equation in the dressed state basis. Transforming into a rotating frame with unitary evolution operator $\mathcal{U}(t)=\exp\left[H_{A}t/i\hbar\right]$, Eq. (27) becomes
@@ IV The Frequency-Filtered Mollow Triplet
The atomic fluorescence is evenly cascaded into an array of $12N+$ 1 tunable single-mode cavities, as depicted in Fig. 7(a). Each mode is modeled as a ring cavity, with two perfect mirrors and two lossy mirrors, each with a loss rate of $\kappa/2$. This results in two separate output channels: a reflection channel consisting of the reflection of photons from, and the transmission of photons out through the input mirror; and a transmission channel consisting of photons emitted out through the second lossy mirror. Each channel is directed towards a separate photodetector (see Fig. 7(b)). The total cascaded system master equation is then [37, 38]
@@ IV.1 Moment equation approach
Figure 7: (a) Full schematic of the multi-mode array-filtered two-level atom, with (b) detail of the decay channels of the $j^{\mathrm{th}}$ mode. The phase-shifted fluorescence is evenly cascaded into a ring cavity with two perfectly reflecting mirrors (bold). The reflection (and transmission) from the input mirror and transmission from the output mirror are sent to separate photo-detectors. (c) The atomic and filter-mode operator moment equations couple in a cascaded scheme, where lower-order moments can be solved independently of higher-order moments. The color of the boxes indicate the order of filter-mode operators, and thus the order in which they can be computed, with the arrows indicating the coupling of lower-order to higher-order moments.

### v3

@@ I Introduction
The majority of studies on atomic resonance fluorescence at this time were largely focused on the entire fluorescence spectrum of the atom. It was at the end of the 1970s when some of the first investigations were conducted on correlations between photons from different components of the Mollow triplet, as opposed to the whole spectrum [9, 10, 11, 12, 13], with the first experimental results reported by Aspect, Roger, Dalibard and Cohen-Tannoudji [14]. To achieve the separation of the two side-peaks, the atomic fluorescence was split into two channels, with each channel passing through an interference filter tuned to the respective target frequency, and modelled as a single-mode optical cavity [15, 16, 17]. Filtering such as this is essential to the measurement of frequency-resoled photon correlations as, typically, photon detectors are board bandwidth, and essentially non-selective in frequency.
With this increased interest in measuring photon correlations of frequency components came interest in the effect that spectral filtering had on measurements [18, 19, 20, 21, 22], as well as interest in other methods of calculating and measuring these correlations. In 2012, del Valle et al. proposed a novel method for measuring frequency-resolved photon correlations [23], consisting of a system of interest weakly coupled to $N$ two-level “detector” atoms, where $N$ is the order of the correlation function of interest. This method has been used extensively in theoretical studies due to its relatively simple implementation [24, 25, 26, 27, 28, 29, 30]. There have been other methods discussed over the years, some with more theoretical leaning, such as: signal processing methods [31]; perturbation approaches [32]; frequency resolved Monte-Carlo, or quantum trajectory methods [33, 34, 35]. Some other methods, such as the eigenvalue decomposition of Liouvillian superoperators developed by Kamide et al [36, 37], allow for the modelling of not only Lorentzian type filters, but for filters with any frequency response.
@@
We also desire that the filter has no effect on the evolution of the source system. This is usually achieved by assuming a vanishingly small coupling of the source system to the filter, such that any back-action can be neglected. Holdaway et al. [32] take this approximation a step further with an algebraic expansion of the source-filter system with respect to the coupling parameter. We can, however, ensure there is no back-action at all by cascading the output of the source system into the filter, using cascaded open quantum systems theory [38, 39]. This can be achieved experimentally through the use of forward propagating waveguides, ring cavities, optical isolators, or beam splitters.
@@
We begin in Section II by introducing the foundations of our filtering model, including a brief discussion of single-mode cavities and rectangular filters. We then present analytic and numerical results for the temporal and frequency response of the multi-mode array filter for different parameters. In Section III we review the dynamics and key phenomena of a resonantly driven two-level atom, which will serve as the source system with which to test the effectiveness of the filter. We then introduce the full atom-filter coupled system in Section IV and the moment equations relevant to solving for the two-time correlation functions. In Section V we present analytic results for frequency-filtered auto- and cross-correlations, comparing the multi-mode array-filtered results with those of a standard Lorentzian filter and approximate ideal solutions derived in the dressed-state basis. Finally, we conclude in Section VI with a summary of key results and a brief discussion on possible experimental implementations.
@@ II.1 Single-mode Lorentzian filter
We begin by considering a standard frequency filtering device, modelled as a single-mode cavity. Applying a coherent driving, with driving strength $\mathcal{E}_{d}$, the Hamiltonian modelling this system is
@@ II.2 The rectangular filter
as shown in Figs. 1(a) and (b). The infinitely sharp frequency response of the sinc, or rectangular, filter, would allow us to maximally increase the bandwidth, and thus the temporal response, of the filter response, much more than a conventional Lorentzian filter, while still maintaining good frequency isolation. Unfortunately, perfect sinc filters are physically impossible to realise due to their non-causal nature, that is, the temporal response depends on future inputs. By neglecting the negative temporal response of the sinc function with $\theta(t)\mathrm{sinc}(at)$, where
@@
is the Heaviside step function, we can realise a causal sinc filter. We see in Fig. 1(d), however, that the frequency response, as calculated from the Fourier transform, changes drastically. The infinitely sharp sides of the rectangular response have been replaced with large slow-decaying wings, together with a large dip in the centre of the response.
Fortunately we are able to recover some of the rectangular nature by introducing a slight phase modulation, or delay, to the temporal response, $\theta(t)\mathrm{sinc}(at-\pi)$ (Fig. 1(e)). With this slight delay, the Fourier transform, Fig. 1(f), shows a much sharper cut-off than the non-delayed response in Fig. 1(d). By increasing the delay even further, we can “recover” much more of the sinc function, resulting in an even more rectangular frequency response. Unfortunately, a greater modulation also results in a greater temporal delay, which is undesirable for measuring accurate photon correlations. We will see in the following section that we are able to recreate this behaviour with a cavity based model.
@@ II.3 The multi-mode array filter
Figure 2: Schematic of a coherently driven multi-mode array filter mode. The input field is evenly split into each individual two-sided cavity mode, where a mode-dependent phase modulation is applied. To achieve this equal coupling of the input light, the input field is passed through an array of beam splitters with increasing intensity reflectivity coefficient. The other (vertical) input field to each beam splitter in the top row is vacuum. Similarly, the input field (horizontal) to the lower output mirror in each filter cavity is also vacuum. Since we consider photon counting measurements, these vacuum fields make no contribution to the detection process.
@@
We then assume that the fields transmitted through the out-coupling mirrors of the cavities are incident on a single photon detector. Note that the input fields to these mirrors are in the vacuum state. As we are measuring the output via photon detectors, these vacuum input fields make no contribution to the detection process. The total electric field at the detector is then given by
@@
where $\omega_{j}$ is the frequency of the $j^{\mathrm{th}}$ cavity mode, $\epsilon_{0}$ is the electric permittivity of free space, and $A$ is the quantisation area of the field. With frequencies in the optical range ($\sim 10^{15}$ Hz), we assume that the effective halfwidth of the multi-mode array filter, $N\delta\omega$, is much smaller than the central mode frequency $\omega_{0}$. The total field may then be approximated as a sum over the annihilation operators:
@@
is the normalised collective mode annihilation operator. For the single-mode case, $N=0$, $\kappa$ is also the halfwidth of the filter response. We therefore define the “effective halfwidth” of the multi-mode array filter as
@@ II.3.1 Temporal response
Figure 3: Normalised field amplitude response to an impulse driving, Eq. (II.3.1), with $m=0$ (blue, solid), $m=1$ (orange, dashed), and $m=2$ (green, dash-dot). Also shown is the sinc term of Eq. (II.3.1) (grey, dotted), showing the effect that the cavity decay, $\kappa$, has on the temporal response. Other parameters are $N=20$ and $\left(\delta\omega,\kappa,\omega_{c}\right)=\left(0.1,0.2,0.0\right)|\mathcal{E}_{d}|.$
@@
We see, then, that this proposed model behaves as we expect from a sinc-response filter, i.e., the temporal response is a positive-sided sinc function, with halfwidth $N\delta\omega$ and a temporal delay of $m\pi$. There is, however, an extra term that arises due to individual cavity modes which make up the multi-mode array filter: the exponential decay and oscillatory term $e^{-\left(\kappa+i\omega_{c}\right)t}$. The oscillatory term, $e^{-i\omega_{c}t}$, simply centres the response in frequency space at $\omega_{c}$. The cavity decay term, however, can have a much more significant effect on the filter’s response. We demonstrate this in Fig. 3, where the temporal response for the phase-shifted multi-mode array (orange, dashed), with $m=0,1$, or 2, is plotted against the sinc term in Eq. (II.3.1). The cavity decay causes the collective temporal response to deviate from the pure sinc term; an effect which is much more noticeable for larger $\kappa$. We can mitigate this effect, and thus achieve a more ideal sinc response, by ensuring the individual cavity decay rate is much smaller than the effective halfwidth of the array filter, i.e., $\kappa\ll N\delta\omega$.
@@ II.3.2 Frequency response
Figure 4: Normalised frequency response for a single-mode cavity with $\kappa=|\mathcal{E}_{d}|$ (grey, dotted) and a multi-mode array filter with (a) no phase modulation $(m=0)$, from Eq. (II.3.2), (b) $\pi$-phase modulation $(m=1)$, and (c) $2\pi$-phase modulation $(m=2)$, with $K=|\mathcal{E}_{d}|$ (blue, solid), $K=2|\mathcal{E}_{d}|$ (orange, dashed), and $K=4|\mathcal{E}_{d}|$ (green, dot-dash). Other parameters are $N=80,\omega_{c}=0$, and, for the multi-mode array filter, $\kappa=0.07|\mathcal{E}_{d}|$.
@@
$|\bar{A}(\omega)|^{2}$ $=\frac{|\mathcal{E}_{d}|^{2}}{\delta\omega^{2}\left(2N+1\right)^{2}}$
@@
We plot the collective frequency response in Fig. 4 for three different sizes of phase modulation – (a) $m=0$, calculated from Eq. (II.3.2), (b) $m=1$ , and (c) $m=2$ – for three different effective halfwidths: $N\delta\omega/|\mathcal{E}_{d}=1$ (blue, solid), 2 (orange, dashed), and 4 (green, dash-dot). By setting $m=0$, the collective response of the multi-mode array filter produces the “bunny-ear” like response, characteristic of the Fourier transform of the positive-sided sinc function in Fig. 1(d). Applying a slight modulation, with $m=1$, we see that the collective response closely resembles that of Fig. 1(f), with much sharper frequency cut-offs than a standard Lorentzian (grey, dotted). We cannot, however, push the phase modulation too far. As previously mentioned, an increase in the size of the modulation also increases the temporal response of the filter. We also see a noticeable dip in the centre of the frequency response when $m=2$, in Fig. 4(c).
@@ III.2 The Mollow triplet
is the normalised steady-state first-order correlation function for the driven two-level atom. In the steady state there exist quantum fluctuations, describable with operators $\Delta\sigma_{\pm}=\sigma_{\pm}-\langle\sigma_{\pm}\rangle_{ss}$, and so we decompose the fluorescence spectrum into coherent and incoherent components:
@@
In the weak driving regime, $\Omega\ll\gamma$, the incoherent power spectrum of the atom is Lorentzian, centred at the atomic frequency $\omega_{A}$, and with halfwidth $\gamma/2$. As the Rabi frequency increases, and thus incoherent scattering dominates, side-peaks emerge at $\omega_{A}\pm\Omega$, giving shape to the Mollow triplet [2, 2], as shown in Fig. 5.
@@ III.3 Atomic dressed states
Each of the Mollow triplet peaks can be attributed to transitions amongst the atomic dressed states [43, 44] – eigenstates of the combined atomic and quantised field Hamiltonians. By diagonalising the Hamiltonian, Eq. (25), we can express the eigenstates in terms of the atomic bare states as:
@@
These dressed states form an energy level ladder, where each pair are separated by an energy of $\hbar\omega_{A}$ in the quantised radiation field. In the strong driving regime, where the two dressed states are well separated, three sets of dressed state transitions occur, as depicted in Fig. 6: |$\ket{d}\rightarrow\ket{u}$ with frequency $\omega_{A}-\Omega$, |$\ket{u}\rightarrow\ket{u}$ and |$\ket{d}\rightarrow\ket{d}$ with frequency $\omega_{A}$, and |$\ket{u}\rightarrow\ket{d}$ with frequency $\omega_{A}+\Omega$. We can express the atomic operators in terms of these dressed states, allowing us to rewrite the atomic master equation in the dressed state basis. Transforming into a rotating frame with unitary evolution operator $\mathcal{U}(t)=\exp\left[H_{A}t/i\hbar\right]$, Eq. (26) becomes
@@ IV The Frequency-Filtered Mollow Triplet
The atomic fluorescence is evenly cascaded into an array of $12N+$ 1 tunable single-mode cavities, as depicted in Fig. 7(a). Each mode is modelled as a ring cavity, with two perfect mirrors and two lossy mirrors, each with a loss rate of $\kappa/2$. This results in two separate output channels: a reflection channel consisting of the reflection of photons from, and the transmission of photons out through the input mirror; and a transmission channel consisting of photons emitted out through the second lossy mirror. Each channel is directed towards a separate photodetector (see Fig. 7(b)). As noted earlier, all input channels to the beam splitters and cavity coupling mirrors that do not emanate from the source are assumed to be vacuum, and thus make no contribution to the detection process. The total cascaded system master equation is then [38, 39]
@@ IV.1 Moment equation approach
Figure 7: (a) Full schematic of the multi-mode array-filtered two-level atom, with (b) detail of the decay channels of the $j^{\mathrm{th}}$ mode. The phase-shifted fluorescence is evenly cascaded into a ring cavity with two perfectly reflecting mirrors (bold). The reflection (and transmission) from the input mirror and transmission from the output mirror are sent to separate photo-detectors. (c) The atomic and filter-mode operator moment equations couple in a cascaded scheme, where lower-order moments can be solved independently of higher-order moments. The colour of the boxes indicate the order of filter-mode operators, and thus the order in which they can be computed, with the arrows indicating the coupling of lower-order to higher-order moments.